### v2

# EVT: Efficient View Transformation for Multi-Modal 3D Object Detection

Abstract
Multi-modal sensor fusion in bird’s-eye-view (BEV) representation has become the leading approach in 3D object detection. However, existing methods often rely on depth estimators or transformer encoders for view transformation, incurring substantial computational overhead. Furthermore, the lack of precise geometric correspondence between 2D and 3D spaces leads to spatial and ray-directional misalignments, restricting the effectiveness of BEV representations. To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency. EVT focuses on two main areas. First, it employs Adaptive Sampling and Adaptive Projection (ASAP), using LiDAR guidance to generate 3D sampling points and adaptive kernels. The generated points and kernels are then used to facilitate the transformation of image features into BEV space and refine the BEV features. Second, EVT includes an improved transformer-based detection framework, which contains a group-wise query initialization method and an enhanced query update framework. It is designed to effectively utilize the obtained multi-modal BEV features within the transformer decoder. By leveraging the geometric properties of object queries, this framework significantly enhances detection performance, especially in a multi-layer transformer decoder structure. EVT achieves state-of-the-art performance on the nuScenes test set with real-time inference speed.
1 Introduction
In the field of 3D object detection, LiDAR-camera fusion is a crucial approach for enhancing perception systems [17, 6, 44, 34, 26, 21, 1, 7, 43, 39, 40, 3]. LiDAR sensors provide precise depth information, which is crucial for accurate object positioning. In contrast, camera sensors capture color and texture, offering rich semantic context. By combining the complementary strengths of these two sensors, robust and reliable 3D object detection can be achieved even in complex environments. This capability is critical in autonomous driving, where consistent and reliable perception is essential.
Figure 1: Performance comparison of EVT and other methods on the nuScenes validation set. The FPS of all methods is measured on a single Tesla A100 GPU using the official implementations, excluding voxelization time.
In multi-modal 3D object detection, the most common approach is to use an ego-centric bird’s-eye-view (BEV) representation space [17, 26, 21, 1, 3]. Early studies employ a 2D-to-3D projection method, which leverages the estimated depth distributions to transform perspective-view images into the BEV space [29, 12, 18, 26, 21]. Although this method utilizes spatial information via pixel-wise depth distributions, its heavy reliance on the performance of the depth estimator limits its overall robustness. Moreover, spatial misalignment frequently occurs when rasterizing the projected 3D points onto the BEV grids.
In contrast, a 3D-to-2D approach [20, 42, 3] has been introduced, where predefined 3D points are projected onto image planes, and features are sampled using deformable attention [49]. However, these predefined points struggle to effectively capture features from regions of interest (ROIs). In addition, the BEV features of the transformed image fail to fully resolve the misalignment in the ray direction.
To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency. EVT focuses on two main areas: (1) adaptive sampling and adaptive projection (ASAP) framework, which is a 3D-to-2D projection-based view transformation method, and (2) an improved transformer-based detection framework, which includes a group-wise query initialization method and an enhanced query update framework for the transformer decoder.
The ASAP framework produces high-quality BEV feature maps from images by leveraging LiDAR guidance, without requiring additional networks, such as depth estimators or transformer encoders. ASAP consists of two main modules. The adaptive sampling (AS) module generates 3D points to transform features from ROIs in images into BEV space. The adaptive projection (AP) module then refines the BEV features obtained from the AS module using adaptive kernels, thereby enhancing feature representation and resolving ray-directional misalignment.
For the improved transformer-based detection framework, the query initialization method uses group-wise learnable parameters, obtaining significant performance improvements in a multi-layer decoder architecture compared to existing methods that initialize queries by sampling features [47, 1]. The enhanced query update framework further refines the deformable cross-attention [49] by integrating corner-aware sampling and position-embedded feature mixing. Corner-aware sampling improves feature sampling in deformable cross-attention by leveraging the geometric properties of object queries. Subsequently, position-embedded feature mixing aggregates the sampled features by explicitly embedding their spatial locations to provide a more detailed representation.
The proposed EVT presents several advantages over existing methods. First, multi-modal fusion in BEV space, without depth estimators or transformer encoders for view transformation, is efficient in terms of computational time and memory usage. Second, the improved transformer-based detection framework that utilizes the characteristics of class groups and geometric properties of objects offers a simple yet powerful solution that can be easily integrated with other transformer-based approaches. As shown in Fig. 1, the proposed EVT achieves 73.7% NDS and 8.3 FPS with ResNet-50 [10], 74.6% NDS and 4.9 FPS with V2-99 [15], and 71.7% NDS and 12.1 FPS with the LiDAR-only model EVT-L on the nuScenes validation set [2], outperforming other methods in both accuracy and inference time. On the nuScenes test set, EVT achieves 75.3% NDS, and 72.5% mAP without model ensemble or test-time augmentation, outperforming previous state-of-the-art methods.
In summary, the main contributions are as follows:
  • The proposed EVT is a novel 3D object detector that leverages a well-structured BEV representation to enhance accuracy and efficiency.
  • The proposed ASAP framework leverages LiDAR guidance to generate high-quality BEV feature maps without substantial computational overhead.
  • The improved transformer-based detection framework enhances overall detection performance and can easily be extended to other transformer-based detection methods.
  • EVT achieves state-of-the-art performance of 75.3% NDS and 72.6% mAP on nuScenes test set.
2 Related Work
Figure 2: Overall architecture of EVT. Each backbone extracts either image features or LiDAR features. The proposed ASAP module fuses these two features in BEV space. The group-wise query initialization stage generates queries from the fused feature map. In the transformer decoder, corner-aware sampling leverages the geometric properties of the object queries to sample multi-modal features, and position-embedded feature mixing decodes the sampled features to update the queries. These queries then predict the 3D bounding boxes.
2.1 Multi-modal Fusion in Transformer
Inspired by the pioneering work of DETR [5], numerous transformer-based detection methods [38, 24, 25, 28, 37, 13, 23, 6, 17, 43, 40] have been proposed. Meanwhile, in multi-modal 3D object detection, extracting features from multiple sensors is still a key challenge. Capturing the complementary relationships between the data from different sensors leads to slow convergence and high computational complexity. To address this issue, FUTR3D [6] proposes a modality-agnostic feature sampler to aggregate features from different sensors using deformable cross attention [49]. DeepInteraction [43] preserves modality-specific information by using a modality interaction strategy. CMT [40] constructs input tokens for the transformer decoder by adding 3D position embeddings to each sensor’s data and then concatenating them. Despite recent advancements, these methods adopt iterative feature fusion in multiple layers, resulting in substantial computational overhead, while the lack of camera parameter utilization limits the effective use of geometric information.
2.2 Multi-modal Fusion in BEV Space
Object detection using a bird’s-eye view (BEV) representation [30, 29, 12, 11, 18, 17, 26, 21, 19, 20, 42, 3] focuses on view transformation of 2D image features into BEV space. For view transformation, 2D-to-3D projection methods map multi-view image features into BEV space by predicting pixel-wise depth distributions [29, 12, 11, 18, 17, 26, 21]. These methods have been extensively studied due to their ability to incorporate spatial priors through depth estimation. BEVFusion [26] fuses LiDAR and camera features in BEV space to provide complementary and robust results. In contrast, 3D-to-2D projection methods project 3D points onto image planes for view transformation without leveraging depth information [19, 20, 42]. BEVFormer [20] projects predefined BEV queries and refines the query features using deformable cross-attention. BEVFusion4D [3] enhances BEV queries by concatenating LiDAR and image BEV features.
However, reliance on inaccurate depth estimation or the use of transformer encoders without fully leveraging geometric properties presents significant challenges in accurately capturing the correlation between 2D and 3D spaces. Furthermore, using additional networks results in substantial computational overhead.
2.3 Query for Transformer-based Detection
Query initialization is crucial in model design for enhancing the performance of transformer-based object detection. Typically, query initialization involves sampling queries based on a heatmap to represent potential object centers. CenterFormer [47] identifies high-scoring keypoints in the heatmap and samples features from the corresponding positions in the input data to initialize queries. TransFusion [1] further incorporates category embeddings into the sampled features for query initialization and fuses these queries with image features in transformer layers. However, this sampling strategy introduces ambiguity by capturing only parts of an object or including multiple object instances, as highlighted by DINO [45].
Furthermore, effectively updating queries within the transformer decoder is essential for enhancing the accuracy of object detection. DETR3D [38] projects 3D learnable queries onto image planes to directly sample the corresponding features. BEVFormer [20] assigns a set of predefined anchor heights to each query on the BEV plane to create 3D reference points and then predicts the 2D sampling offsets on the image planes after projection. SparseBEV [23] generates 3D sampling offsets to formulate 3D sampling points by applying bounding box size and rotation to each query, and compensates sampling points based on the object and ego motion across multiple frames. Although existing methods have made significant progress, there remains considerable potential to further leverage geometric properties for 3D object detection.
3 Methodology
The proposed EVT is a novel method for 3D object detection with effective multi-modal fusion and improved transformer-based detection framework. The overall pipeline is shown in Fig. 2. First, $N_{\text{s}}$-scale perspective-view image features $\{\mathbf{PV}_{j}\mid j\in\{1,2,\dots,N_{\text{s}}\}\}$ and bird’s-eye-view (BEV) LiDAR features $\mathbf{BEV}_{\text{lidar}}\in\mathbb{R}^{C\times H\times W}$ are extracted from individual backbones, where $C$ is the feature dimension, and $H\times W$ is the size of the BEV feature map. Next, the proposed adaptive sampling and adaptive projection (ASAP) module fuses the multi-modal features in BEV space (Sec. 3.1). Group-wise queries are then initialized in BEV space using the heatmap head (Sec. 3.2), and updated in the transformer decoder based on the geometric properties of the 3D bounding box (Sec. 3.3).
3.1 Adaptive Sampling and Adaptive Projection
The proposed ASAP module effectively leverages LiDAR guidance to transform multi-scale perspective-view image features into BEV space. The structure of this module is shown in Fig. 3.
Adaptive Sampling
The adaptive sampling (AS) module samples multi-scale image features into BEV space by predicting the height of each sampling point on the BEV plane and assigning weights to the predicted heights and image scales.
First, a set of heights for each grid cell is generated from LiDAR features:
$\{Z_{i}\}_{i=1}^{N_{\text{h}}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v),$ (1)
where $(u,v)$ denotes the coordinates of a grid cell in BEV space, and $N_{\text{h}}$ denotes the number of generated heights. Consequently, the 3D adaptive sampling points $P=\{(X,Y,Z_{i})\}_{i=1}^{N_{\text{h}}}$ are defined using real-world coordinates $(X,Y)$ corresponding to the grid cell at $(u,v)$ and the set of heights $\{Z_{i}\}$.
Next, the generated points $P$ are projected onto the image plane. Each projected point samples the multi-scale features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ across $N_{\text{s}}$ different image scales with downsampled strides $\{S_{j}\}_{j=1}^{N_{\text{s}}}$. Consequently, $N_{\text{h}}\times N_{\text{s}}$ projected points $(x^{j}_{i},y^{j}_{i})$ and sampled features $f^{j}_{i}$ are obtained from the projection:
$(x^{j}_{i},y^{j}_{i})=\frac{{\rm Proj}(X,Y,Z_{i})}{S_{j}}$ (2)
$f^{j}_{i}={\rm B}(\mathbf{PV}_{j},(x^{j}_{i},y^{j}_{i}))\in\mathbb{R}^{C},$ (3)
where ${\rm Proj}(\cdot)$ denotes the projection of 3D points onto the image plane, and ${\rm B}(\cdot)$ denotes bilinear interpolation.
To aggregate all the $N_{\text{s}}\times N_{\text{h}}$ sampled features $\{f^{j}_{i}\}$, adaptive sampling weights $W_{\text{as}}\in\mathbb{R}^{N_{\text{s}}\times N_{\text{h}}}$ are obtained from $\mathbf{BEV}_{\text{lidar}}$. $W_{\text{as}}$ determines the importance of the heights and image feature scales for each grid cell. The multi-scale image features in BEV space are obtained as follows:
$W_{\text{as}}=\sigma({\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v))$ (4)
$\mathbf{BEV}_{\text{as}}(u,v)=\sum_{j=1}^{N_{\text{s}}}\sum_{i=1}^{N_{\text{h}}}W_{\text{as}}(j,i)\cdot f^{j}_{i},$ (5)
where $\sigma(\cdot)$ denotes the softmax function applied to all $N_{\text{s}}\times N_{\text{h}}$ elements. $\mathbf{BEV}_{\text{as}}$ denotes the image BEV feature map, representing the features of the ROIs from the images in BEV space by leveraging LiDAR guidance.
Figure 3: Overview of AS and AP. In AS, the LiDAR BEV feature map generates 3D sampling points and their corresponding weights. The sampling points are projected onto the image plane, where features are sampled and combined by weighted pooling in each BEV grid cell. In AP, the image BEV feature map produced by AS is refined channel-wise using the adaptive kernels generated from the LiDAR features.
Adaptive Projection
The adaptive projection (AP) module refines the BEV feature $\mathbf{BEV}_{\text{as}}$ produced by the AS module using adaptive kernels for each grid cell. The overall process is represented by the following equations:
$K_{\text{ap}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v)\in\mathbb{R}^{C\times C}$ (6)
$\mathbf{BEV}_{\text{camera}}(u,v)=\mathbf{BEV}_{\text{as}}(u,v)\times K_{\text{ap}}\in\mathbb{R}^{C}.$ (7)
First, an adaptive kernel $K_{\text{ap}}\in\mathbb{R}^{C\times C}$ is generated from $\mathbf{BEV}_{\text{lidar}}$. This kernel is generated for each BEV grid cell to enable detailed feature representation. Subsequently, $\mathbf{BEV}_{\text{as}}$ is refined by applying a channel-wise linear projection using $K_{\text{ap}}$ to obtain the image BEV feature map $\mathbf{BEV}_{\text{camera}}$.
This LiDAR-guided refinement improves feature representation and addresses the ray-directional misalignment inherent in 3D-to-2D projection methods, which suffer from occlusions or empty 3D spaces.
Multi-modal Fusion in BEV Space
The multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}\in\mathbb{R}^{C\times H\times W}$ is obtained by channel-wise concatenation of the image and LiDAR BEV feature maps, followed by a convolution operation:
$\mathbf{BEV}_{\text{fuse}}={\rm Conv}({\rm Concat}(\mathbf{BEV}_{\text{camera}},\mathbf{BEV}_{\text{lidar}})),$ (8)
where ${\rm Concat}(\cdot)$ denotes the channel-wise concatenation.
Method Modality NDS (val) mAP (val) NDS (test) mAP (test)
UVTR-L [17] L 67.7 60.9 69.7 63.9
TransFusion-L [1] L 70.1 65.1 70.2 65.5
FocalFormer3D-L [7] L - - 72.6 68.7
CMT-L [40] L 68.6 62.4 70.1 65.3
EVT-L (Ours) L 71.7 66.4 72.1 67.7
MVP [44] LC 70.8 67.1 70.5 66.4
UVTR [17] LC 70.2 65.4 71.1 67.1
AutoAlignV2 [8] LC 71.2 67.1 72.4 68.4
TransFusion [1] LC 71.3 67.5 71.7 68.9
DeepInteraction [43] LC 72.6 69.9 73.4 70.8
BEVFusion [26] LC 71.4 68.5 72.9 70.2
Objectfusion [4] LC 72.3 69.8 73.3 71.0
FocalFormer3D [7] LC 71.1 66.5 73.9 71.6
CMT [40] LC 72.9 70.3 74.1 72.0
BEVFusion4D-S [3] LC 72.9 70.9 73.7 71.9
SparseFusion [39] LC 72.8 70.4 73.8 72.0
MSMDFusion [14] LC - - 74.0 71.5
UniTR [35] LC 73.3 70.5 74.5 70.9
EVT (Ours) LC 74.6 72.1 75.3 72.6
Table 1: Performance comparison on the nuScenes validation and test sets. The results are obtained without model ensemble or test-time augmentation. ‘L’ and ‘C’ denote LiDAR and Camera, respectively.
3.2 Group-wise Query Initialization
The proposed group-wise query initialization method effectively generates queries for transformer-based detection frameworks. First, group-wise heatmaps are predicted from the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$. The groups are predefined, with each group containing similarly-sized classes. The predicted heatmaps have scores ranging from 0 to 1, representing the likelihood that each BEV pixel corresponds to the center of an object. The heatmap head is supervised by 2D Gaussian distributions centered at each object’s location. Next, the top-k keypoints are selected from each heatmap group. The positions of these keypoints are used as reference points for queries in BEV space.
This method initializes each query feature using group-wise learnable parameters. Queries within the same group share these parameters, which enables them to effectively capture the characteristics of group. This demonstrates better performance in the multi-layer transformer decoder than traditional query initialization approaches that sample features directly from the input feature map [47, 1]. Further details are provided in Sec. 4.4.
3.3 Enhanced Query Update Framework
For more effective query updates within a transformer decoder based on deformable attention [49], we propose the corner-aware sampling and position-embedded feature mixing methods.
Corner-aware Sampling.
The proposed corner-aware sampling method shifts the sampling reference from the center of the object to each corner and rotates sampling offsets according to the object’s heading. This approach addresses the limitations of existing center-based sampling methods, which generate sampling offsets around the object’s center, resulting in difficulties capturing the full spatial extent and shape of the objects.
First, the initial points $\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\}$ are generated from the query feature $\mathbf{q}$ using a linear layer:
$\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\mid i\in 0,1,...,N_{\text{p}}-1\}={\rm Linear}(\mathbf{q}),$ (9)
where $N_{\text{p}}$ denotes the number of sampling points.
Subsequently, the sampling offsets $\{(\Delta x_{i},\Delta y_{i})\}$ and sampling points $\{(x_{i},y_{i})\}$ are determined via a geometric transformation as follows:
$\left[\!\!\begin{array}[]{c}\Delta x_{i}\\[5.0pt] \Delta y_{i}\\ \end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}\cos\theta\!\!\!&-\!\sin\theta\!\!\\[5.0pt] \sin\theta\!\!\!&\cos\theta\!\!\!\!\\ \end{array}\!\!\right]\!\!\left[\!\!\begin{array}[]{c}{\rm I}_{j}\cdot\frac{l}{2}+\Delta x_{i}^{\prime}\\[5.0pt] {\rm I}_{j}^{\prime}\cdot\!\frac{w}{2}\!+\Delta y_{i}^{\prime}\\ \end{array}\!\!\right]\!$ (10)
$\left[\!\!\begin{array}[]{cc}x_{i}&y_{i}\!\end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}x_{c}&y_{c}\end{array}\!\!\right]\!+\!\left[\!\begin{array}[]{cc}\!\Delta x_{i}&\Delta y_{i}\!\!\\ \end{array}\!\right],$ (11)
where $(I_{j},I^{\prime}_{j})\,\in\{(1,1),(1,-1),(-1,1),(-1,-1)\}$. $j$ denotes the index of each corner and is the remainder of $i$ divided by four, the number of corners. $(x_{c},y_{c})$ denote the coordinates of the object’s center. $l$, $w$, and $\theta$ denote the length, width, and yaw of the predicted bounding box, respectively, obtained from the regression head of the previous transformer decoder layer. In the first decoder layer, where there is no preceding layer to provide bounding boxes, the equation is satisfied by setting $l$, $w$, and $\theta$ to zero.
The features are sampled by bilinear interpolation at each sampling point on the multi-modal fused features $\mathbf{BEV}_{\text{fuse}}$ as follows:
$g_{i}={\rm B}(\mathbf{BEV}_{\text{fuse}},(x_{i},y_{i}))\in\mathbb{R}^{C},$ (12)
where ${\rm B}(\cdot)$ denotes the bilinear interpolation. $g_{i}$ represents the sampled feature that captures important key tokens by leveraging the geometric properties of the query’s bounding box.
Position-embedded Feature Mixing.
Accurate locations of the sampled features are essential for providing a more detailed representation of the bounding boxes. In AdaMixer [9], the adaptive channel and spatial mixing are combined to leverage query guidance and effectively decode sampled features. However, this method is less effective for decoding the outputs of the corner-aware sampling, as shown in Tab. 3 (e), which involves complex geometric transformations of the sampling points. To address this, we propose position-embedded feature mixing, an enhancement based on AdaMixer that improves the decoding capability by incorporating embeddings for sampling offsets.
First, the sampling offsets are embedded as position vectors using sinusoidal position encoding [33], followed by a linear layer. Then, the position-embedded sampled features $G\in\mathbb{R}^{N_{\text{p}}\times C}$ are obtained by element-wise addition of the sampled feature $g_{i}$ and the position vector $e_{i}$:
$e_{i}$ $={\rm Linear}(\Phi_{pos}((x_{i},y_{i})))\in\mathbb{R}^{C}$ (13)
$G$ =$=\{(g_{i}+e_{i})\mid i\in{0,1,...,N_{\text{p}}-1}\},$ (14)
where $\Phi_{pos}(\cdot)$ denotes sinusoidal position encoding.
Subsequently, adaptive channel mixing is applied to $G$ using the dynamic weights $W_{c}$ generated from the query feature $\mathbf{q}$ to obtain the channel-mixed feature $G_{c}$:
$W_{c}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{C\times C}$ (15)
$G_{c}$ $={\rm ReLU}({\rm LN}(G\times W_{c})).$ (16)
Next, adaptive spatial mixing is applied to the spatial dimensions of $G_{c}$ using dynamic weights $W_{s}$ to obtain the spatial-mixed feature $G_{cs}$:
$W_{s}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{N_{\text{p}}\times N_{\text{p}}}$ (17)
$G_{cs}$ $={\rm ReLU}({\rm LN}(G_{c}^{T}\times W_{s})).$ (18)
Finally, the query feature is updated as follows:
$\mathbf{q^{\prime}}=\mathbf{q}+{\rm Linear}({\rm Flatten}(G_{cs})),$ (19)
where $\mathbf{q^{\prime}}$ represents the updated query feature that is enhanced for a more accurate feature representation by incorporating the positional information of each sampled point.
4 Experiments
Figure 4: Visualization of projected sampling points for each object. The top row shows the projections of predefined 3D points, and the bottom row shows the projections of the points generated by the AS module. Points of each object are denoted by different colors.
Figure 5: Comparison of BEV feature maps from different methods: (a) Vanilla, (b) AS, (c) AP, and (d) ASAP. Vanilla shows less informative and unaligned features. AS improves feature representation using adaptive sampling. AP corrects ray-directional misalignment. ASAP integrates both AS and AP, leading to the most refined and well-aligned feature representation.
4.1 Implementation Details
For the image backbone, either ResNet-50 [10] with a resolution of $704\times 256$ or V2-99 [15] with $1600\times 640$ is used with the feature pyramid network (FPN) [22]. For the LiDAR backbone, VoxelNet [46] is used with an ROI of $[-54.0m,54.0m]$ along $(X,Y)$ and $[-5.0m,3.0m]$ along $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$. In ASAP, we use 4 points for each grid cell and the last two stages of each image backbone for view transformation. The size of the unified BEV feature map size is $180\times 180$. The object categories are divided into 6 groups, with 150 queries per group, resulting in a total of 900 queries. The transformer decoder has 6 layers, and the corner-aware sampling generates 16 sampling points per query. The feature dimension is set to 256.
Our model is trained on 8 RTX 3090 GPUs with a batch size of 16. The model is trained end-to-end for 10 epochs using CBGS [48], whereas GT sample augmentation [41] is applied for the first 9 epochs. The query denoising strategy [16] is also adopted. Gaussian Focal loss [36], Focal loss [31] and L1 loss are used for heatmap prediction, classification and regression, respectively. The AdamW [27] optimizer is adopted with a learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-2}$ with a cyclical learning rate policy [32]. No model ensemble or test-time augmentation is applied during inference.
4.2 Dataset and Metric
[…]
Performance on this dataset is assessed using metrics like mean Average Precision (mAP), calculated over distance thresholds of 0.5, 1, 2, and 4m across all classes, and the nuScenes Detection Score (NDS), which offers a holistic evaluation by combining mAP with measures of translation, scale, orientation, velocity, and attribute errors.
4.3 Comparison with State-of-the-art Methods
As shown in Tab. 1, we compare EVT and its LiDAR-only model, EVT-L, with existing methods on the nuScenes validation and test sets. The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, outperforming all existing state-of-the-art approaches on both the nuScenes validation and test sets. In particular, it surpasses recent methods, such as UniTR [35] by 0.8% NDS and 1.7% mAP, MSMDFusion [14] by 1.3% NDS and 1.1% mAP, and SparseFusion [39] by 1.5% NDS and 0.6% mAP. Furthermore, EVT shows a performance improvement of 3.2% NDS and 4.9% mAP compared with the LiDAR-only model EVT-L. In contrast, TransFusion [1] shows only a 1.5% NDS and 3.4% mAP improvement over its LiDAR-only model, TransFusion-L. This indicates that EVT effectively utilizes camera data through the proposed view transformation method.
We also compare EVT-L with the LiDAR-only versions of other multi-modal 3D object detectors. On the nuScenes validation set, EVT-L surpasses TransFusion [1] and CMT [40] by 1.6% and 3.1% NDS, respectively. EVT-L also demonstrates competitive performance on the test set. These results validate the effectiveness of the proposed query initialization and cross-attention mechanisms in enhancing 3D object detection.
LiDAR Camera AS AP NDS mAP
(a) 71.7 66.4
(b) 72.7 69.1
(c) 73.5 70.6
(d) 73.7 70.7
Table 2: Ablation study of the proposed VT method. The proposed ASAP shows a significant performance improvement compared to the vanilla VT method (b), which projects predefined 3D points.
Figure 6: Comparison of query initialization strategies on the nuScenes validation set. For the multi-layer transformer decoder, group-wise query initialization without feature sampling achieves the best performance.
Attention Reference scale rotate feature mixing position-embedded NDS mAP
(a) Standard Center 69.6 65.3
(b) Deformable Center 70.0 64.8
(c) 70.5 65.2
(d) Ours Corner 71.3 65.7
(e) 71.2 65.6
(f) 71.7 66.4
Table 3: Comparison of the cross-attention methods within the transformer decoder. The proposed corner-aware sampling and position-embedded feature mixing achieve significant performance improvements.
Initial Query Formulation NDS mAP
BEV Reference Point 71.7 66.4
Bbox from regression head 71.4 66.3
Bbox from learnable parameters 70.9 66.0
Table 4: Comparison of the initial query formulations for the first layer of the transformer decoder.
4.4 Ablation Studies
In this section, we describe the validation of each component of the proposed method on the nuScenes validation set. Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [48] and the query denoising strategy [16].
Adaptive Sampling and Adaptive Projection
Tab. 2 displays the results of the ablation conducted on ASAP by retraining the entire model. In these experiments, ResNet [10] is used as the backbone network with a resolution of $704\times 256$. (a) shows the performance of EVT-L, the LiDAR-only model. In (b), the vanilla view transformation (VT) method, without LiDAR guidance, employs a 3D-to-2D projection of predefined 3D points for feature sampling.
In (c), AS achieves improvements of 0.8% NDS and 1.5% mAP compared to the vanilla VT method. The sampling points of the vanilla VT and AS are visualized in Fig. 4. The sampling points of AS are adaptively generated in the ROIs.
In (d), the entire ASAP achieves improvements of 1.0% NDS and 1.6% mAP compared to the vanilla VT method (b). The BEV feature maps of each component are visualized in Fig. 5. ASAP effectively transforms image features into BEV space and resolves ray-directional misalignment.
Group-wise Query Initialization
Different query initialization methods using the heatmap head are ablated in Fig. 6. All methods utilize a top-K keypoints selection approach to define queries, but differ in their specific implementations. (a) directly samples features from the input BEV feature map and combines them with group-wise embedding, which is similar to the approach commonly used in 3D object detectors [47, 1]. (b) initializes queries with group-wise embedding using global top-K selection. (c) is our proposed group-wise query initialization strategy, which initializes queries using top-K selection and embedding for each group.
In the results, the direct feature sampling approach in (a) is more effective in a single-layer transformer decoder, outperforming (c) by 0.6% NDS. However, in a multi-layer transformer decoder, this leads to performance degradation. Specifically, with 6 layers, (c) surpasses (a) by 1.0% NDS. Additionally, group-wise top-K (c) consistently outperforms global top-K (b) across all layers by approximately 0.1% NDS.
Enhanced Query Update Framework
We ablate the corner-aware sampling method, as demonstrated in Tab. 3 (a)-(d). (a) is the standard multi-head cross attention [33], and (b)-(c) are the standard deformable attention [49]. Specifically, in (c), the sampling offsets are scaled according to the bounding box size and rotated according to the heading. (d) is the proposed corner-aware sampling method, which samples features from the bounding box corners of each query and rotates them according to the heading. The corner-aware sampling method (d) achieves improvements of 0.8% NDS and 0.5% mAP compared to method (c).
The results of the ablation study on the feature mixing method are displayed in Tab. 3 (e)-(f). In (e), the standard feature mixing method of AdaMixer [9] does not yield any improvements over our method. In (f), the proposed position-embedded feature mixing demonstrates improvements of 0.4% NDS and 0.7% mAP compared to (d).
Initial Query Formulation for Transformer
As described in Sec. 3.3, the corner-aware sampling is applied starting from the second transformer decoder layer. We experimented on two different approaches to incorporate the corner-aware sampling into all transformer decoder layers. The first approach involves adding regression heads during the query initialization stage (Sec. 3.2) to predict the bounding boxes, whereas the second approach uses learnable bounding boxes from SparseBEV [23]. However, as shown in Tab. 4, neither of these approaches led to any noticeable performance improvement. Furthermore, the experiment using learnable bounding boxes results in performance degradation.
5 Conclusion
In this paper, we have proposed EVT, a novel transformer-based 3D object detector. The proposed view transformation method, ASAP, enables robust and efficient LiDAR-camera fusion by leveraging LiDAR guidance. In addition, the proposed group-wise query initialization method provides effective queries for a multi-layer transformer decoder. The corner-aware sampling and position-embedded feature mixing methods effectively utilize geometric properties to update queries, and these methods can easily be extended to other models. We expect EVT to provide valuable insights into multi-modal 3D object detection.

### v3

@@
 Abstract
-Multi-modal sensor fusion in bird’s-eye-view (BEV) representation has become the leading approach in 3D object detection.
-However, existing methods often rely on depth estimators or transformer encoders for view transformation, incurring substantial computational overhead.
-Furthermore, the lack of precise geometric correspondence between 2D and 3D spaces leads to spatial and ray-directional misalignments, restricting the effectiveness of BEV representations.
-To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency.
-EVT focuses on two main areas.
-First, it employs Adaptive Sampling and Adaptive Projection (ASAP), using LiDAR guidance to generate 3D sampling points and adaptive kernels.
-The generated points and kernels are then used to facilitate the transformation of image features into BEV space and refine the BEV features.
-Second, EVT includes an improved transformer-based detection framework, which contains a group-wise query initialization method and an enhanced query update framework.
-It is designed to effectively utilize the obtained multi-modal BEV features within the transformer decoder.
-By leveraging the geometric properties of object queries, this framework significantly enhances detection performance, especially in a multi-layer transformer decoder structure.
-EVT achieves state-of-the-art performance on the nuScenes test set with real-time inference speed.
+Multi-modal sensor fusion in Bird’s Eye View (BEV) representation has become the leading approach for 3D object detection.
+However, existing methods often rely on depth estimators or transformer encoders to transform image features into BEV space, which reduces robustness or introduces significant computational overhead.
+Moreover, the insufficient geometric guidance in view transformation results in ray-directional misalignments, limiting the effectiveness of BEV representations.
+To address these challenges, we propose Efficient View Transformation (EVT), a novel 3D object detection framework that constructs a well-structured BEV representation, improving both accuracy and efficiency.
+Our approach focuses on two key aspects.
+First, Adaptive Sampling and Adaptive Projection (ASAP), which utilizes LiDAR guidance to generate 3D sampling points and adaptive kernels, enables more effective transformation of image features into BEV space and a refined BEV representation.
+Second, an improved query-based detection framework, incorporating group-wise mixed query selection and geometry-aware cross-attention, effectively captures both the common properties and the geometric structure of objects in the transformer decoder.
+On the nuScenes test set, EVT achieves state-of-the-art performance of 75.3% NDS with real-time inference speed.
 1 Introduction
-In the field of 3D object detection, LiDAR-camera fusion is a crucial approach for enhancing perception systems [17, 6, 44, 34, 26, 21, 1, 7, 43, 39, 40, 3].
-LiDAR sensors provide precise depth information, which is crucial for accurate object positioning.
-In contrast, camera sensors capture color and texture, offering rich semantic context.
-By combining the complementary strengths of these two sensors, robust and reliable 3D object detection can be achieved even in complex environments.
-This capability is critical in autonomous driving, where consistent and reliable perception is essential.
+LiDAR-camera fusion is essential for 3D object detection, as it leverages the complementary strengths [21, 5, 53, 39, 31, 25, 1, 6, 49, 45, 46, 3].
+LiDAR provides precise geometric information for accurate object localization, while cameras capture rich semantic details such as color and texture.
+However, their integration remains challenging due to the differences in sensing modalities.
 Figure 1: Performance comparison of EVT and other methods on the nuScenes validation set.
-The FPS of all methods is measured on a single Tesla A100 GPU using the official implementations, excluding voxelization time.
-In multi-modal 3D object detection, the most common approach is to use an ego-centric bird’s-eye-view (BEV) representation space [17, 26, 21, 1, 3].
-Early studies employ a 2D-to-3D projection method, which leverages the estimated depth distributions to transform perspective-view images into the BEV space [29, 12, 18, 26, 21].
-Although this method utilizes spatial information via pixel-wise depth distributions, its heavy reliance on the performance of the depth estimator limits its overall robustness.
-Moreover, spatial misalignment frequently occurs when rasterizing the projected 3D points onto the BEV grids.
-In contrast, a 3D-to-2D approach [20, 42, 3] has been introduced, where predefined 3D points are projected onto image planes, and features are sampled using deformable attention [49].
-However, these predefined points struggle to effectively capture features from regions of interest (ROIs).
-In addition, the BEV features of the transformed image fail to fully resolve the misalignment in the ray direction.
-To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency.
-EVT focuses on two main areas: (1) adaptive sampling and adaptive projection (ASAP) framework, which is a 3D-to-2D projection-based view transformation method, and (2) an improved transformer-based detection framework, which includes a group-wise query initialization method and an enhanced query update framework for the transformer decoder.
-The ASAP framework produces high-quality BEV feature maps from images by leveraging LiDAR guidance, without requiring additional networks, such as depth estimators or transformer encoders.
-ASAP consists of two main modules.
-The adaptive sampling (AS) module generates 3D points to transform features from ROIs in images into BEV space.
-The adaptive projection (AP) module then refines the BEV features obtained from the AS module using adaptive kernels, thereby enhancing feature representation and resolving ray-directional misalignment.
-For the improved transformer-based detection framework, the query initialization method uses group-wise learnable parameters, obtaining significant performance improvements in a multi-layer decoder architecture compared to existing methods that initialize queries by sampling features [47, 1].
-The enhanced query update framework further refines the deformable cross-attention [49] by integrating corner-aware sampling and position-embedded feature mixing.
-Corner-aware sampling improves feature sampling in deformable cross-attention by leveraging the geometric properties of object queries.
-Subsequently, position-embedded feature mixing aggregates the sampled features by explicitly embedding their spatial locations to provide a more detailed representation.
-The proposed EVT presents several advantages over existing methods.
-First, multi-modal fusion in BEV space, without depth estimators or transformer encoders for view transformation, is efficient in terms of computational time and memory usage.
-Second, the improved transformer-based detection framework that utilizes the characteristics of class groups and geometric properties of objects offers a simple yet powerful solution that can be easily integrated with other transformer-based approaches.
+The FPS of all methods is measured in FP32 on a single Tesla A100 GPU using the official implementations, excluding voxelization time.
+Recently, the dominant multi-modal fusion methods are categorized as implicit or explicit fusion.
+Implicit fusion employs cross-attention within transformers, where object queries iteratively interact with independently processed sensor features [5, 1, 49, 50, 46].
+Implicit fusion offers flexibility and simplicity by removing explicit feature alignment across sensor modalities but incurs high computational costs and struggles to extract complementary features.
+Explicit fusion, on the other hand, directly aligns and integrates sensor data in BEV space using view transformation (VT), which enhances complementary feature fusion while reducing computational costs.
+Although its performance heavily depends on the accuracy of VT, conventional methods have inherent limitations.
+Depth-based VT methods [34, 22, 31, 25, 56, 20] lift image features into BEV using pixel-wise depth estimation, but their sensitivity to depth errors compromises robustness.
+In contrast, query-based VT methods [24, 1, 8, 7, 3, 12] refine BEV queries via attention mechanisms, incurring high computational costs.
+Furthermore, both methods lack geometric guidance, resulting in ray-directional misalignment.
+This misalignment reduces spatial accuracy, leading to unintended information capture along the ray direction, ultimately degrading BEV representation accuracy.
+To address the limitations of explicit fusion methods and improve detection performance, we propose Efficient View Transformation (EVT), a novel 3D object detection framework aimed at enhancing both accuracy and efficiency by constructing a well-structured BEV representation.
+EVT introduces two key innovations: (1) Adaptive Sampling and Adaptive Projection (ASAP), a novel VT method which leverages LiDAR guidance, and (2) an improved query-based object detection framework incorporating group-wise mixed query selection and geometry-aware cross-attention, enabling effective multi-modal BEV feature decoding for accurate 3D object detection.
+ASAP consists of two key modules: Adaptive Sampling (AS) and Adaptive Projection (AP).
+AS generates 3D sampling points from LiDAR features to effectively represent image features in BEV space while focusing more on high-relevance areas in the image.
+AP refines BEV representations using adaptive kernels generated from LiDAR features to enhance structurally meaningful 3D information.
+As a result, ASAP improves BEV feature representation and eliminates ray-directional misalignment.
+Unlike existing methods, it does not rely on depth estimation or attention mechanism, ensuring both efficiency and robustness.
+Additionally, we improve the query-based object detection framework by introducing group-wise mixed query selection and geometry-aware cross-attention.
+The mixed query selection generates object queries using group-wise learnable parameters and heatmaps, allowing them to capture the common properties of each group and initialize at high-confidence positions.
+Then, the geometry-aware cross-attention refines object queries by integrating corner-aware sampling for precise feature selection and position-aware feature mixing for spatially aware feature decoding.
+These enhancements improve more robust and accurate detection while maintaining computational efficiency.
+We evaluate EVT on the nuScenes dataset in terms of accuracy and efficiency.
 As shown in Fig.
-1, the proposed EVT achieves 73.7% NDS and 8.3 FPS with ResNet-50 [10], 74.6% NDS and 4.9 FPS with V2-99 [15], and 71.7% NDS and 12.1 FPS with the LiDAR-only model EVT-L on the nuScenes validation set [2], outperforming other methods in both accuracy and inference time.
-On the nuScenes test set, EVT achieves 75.3% NDS, and 72.5% mAP without model ensemble or test-time augmentation, outperforming previous state-of-the-art methods.
+1, EVT achieves 74.1% NDS and 8.3 FPS with ResNet-50 [11], 74.6% NDS and 4.9 FPS with V2-99 [18], and 71.7% NDS and 12.1 FPS with the LiDAR-only model EVT-L on the nuScenes validation set, outperforming other methods in both accuracy and inference speed.
+On the nuScenes test set, EVT achieves 75.3% NDS and 72.5% mAP using single-frame raw data, without model ensemble or test-time augmentation, surpassing previous state-of-the-art methods.
 In summary, the main contributions are as follows:
-  • The proposed EVT is a novel 3D object detector that leverages a well-structured BEV representation to enhance accuracy and efficiency.
-  • The proposed ASAP framework leverages LiDAR guidance to generate high-quality BEV feature maps without substantial computational overhead.
-  • The improved transformer-based detection framework enhances overall detection performance and can easily be extended to other transformer-based detection methods.
-  • EVT achieves state-of-the-art performance of 75.3% NDS and 72.6% mAP on nuScenes test set.
+  • We introduce EVT, a novel 3D object detection framework that improves both accuracy and efficiency through ASAP and an improved query-based framework.
+  • ASAP, our view transformation method, utilizes LiDAR guidance to generate BEV feature maps while improving efficiency by eliminating the need for depth estimators and transformer encoders.
+  • Our improved query-based detection framework further enhances detection accuracy and robustness.
+  • EVT achieves state-of-the-art performance of 75.3% NDS and 72.6% mAP on the nuScenes test set.
 2 Related Work
@@
 The proposed ASAP module fuses these two features in BEV space.
-The group-wise query initialization stage generates queries from the fused feature map.
-In the transformer decoder, corner-aware sampling leverages the geometric properties of the object queries to sample multi-modal features, and position-embedded feature mixing decodes the sampled features to update the queries.
+The group-wise mixed query selection stage generates queries from the fused feature map.
+In the transformer decoder, corner-aware sampling leverages the geometric properties of the object queries to sample multi-modal features, and position-aware feature mixing decodes the sampled features to update the queries.
 These queries then predict the 3D bounding boxes.
-2.1 Multi-modal Fusion in Transformer
-Inspired by the pioneering work of DETR [5], numerous transformer-based detection methods [38, 24, 25, 28, 37, 13, 23, 6, 17, 43, 40] have been proposed.
-Meanwhile, in multi-modal 3D object detection, extracting features from multiple sensors is still a key challenge.
-Capturing the complementary relationships between the data from different sensors leads to slow convergence and high computational complexity.
-To address this issue, FUTR3D [6] proposes a modality-agnostic feature sampler to aggregate features from different sensors using deformable cross attention [49].
-DeepInteraction [43] preserves modality-specific information by using a modality interaction strategy.
-CMT [40] constructs input tokens for the transformer decoder by adding 3D position embeddings to each sensor’s data and then concatenating them.
-Despite recent advancements, these methods adopt iterative feature fusion in multiple layers, resulting in substantial computational overhead, while the lack of camera parameter utilization limits the effective use of geometric information.
-2.2 Multi-modal Fusion in BEV Space
-Object detection using a bird’s-eye view (BEV) representation [30, 29, 12, 11, 18, 17, 26, 21, 19, 20, 42, 3] focuses on view transformation of 2D image features into BEV space.
-For view transformation, 2D-to-3D projection methods map multi-view image features into BEV space by predicting pixel-wise depth distributions [29, 12, 11, 18, 17, 26, 21].
-These methods have been extensively studied due to their ability to incorporate spatial priors through depth estimation.
-BEVFusion [26] fuses LiDAR and camera features in BEV space to provide complementary and robust results.
-In contrast, 3D-to-2D projection methods project 3D points onto image planes for view transformation without leveraging depth information [19, 20, 42].
-BEVFormer [20] projects predefined BEV queries and refines the query features using deformable cross-attention.
-BEVFusion4D [3] enhances BEV queries by concatenating LiDAR and image BEV features.
-However, reliance on inaccurate depth estimation or the use of transformer encoders without fully leveraging geometric properties presents significant challenges in accurately capturing the correlation between 2D and 3D spaces.
-Furthermore, using additional networks results in substantial computational overhead.
-2.3 Query for Transformer-based Detection
-Query initialization is crucial in model design for enhancing the performance of transformer-based object detection.
-Typically, query initialization involves sampling queries based on a heatmap to represent potential object centers.
-CenterFormer [47] identifies high-scoring keypoints in the heatmap and samples features from the corresponding positions in the input data to initialize queries.
-TransFusion [1] further incorporates category embeddings into the sampled features for query initialization and fuses these queries with image features in transformer layers.
-However, this sampling strategy introduces ambiguity by capturing only parts of an object or including multiple object instances, as highlighted by DINO [45].
-Furthermore, effectively updating queries within the transformer decoder is essential for enhancing the accuracy of object detection.
-DETR3D [38] projects 3D learnable queries onto image planes to directly sample the corresponding features.
-BEVFormer [20] assigns a set of predefined anchor heights to each query on the BEV plane to create 3D reference points and then predicts the 2D sampling offsets on the image planes after projection.
-SparseBEV [23] generates 3D sampling offsets to formulate 3D sampling points by applying bounding box size and rotation to each query, and compensates sampling points based on the object and ego motion across multiple frames.
-Although existing methods have made significant progress, there remains considerable potential to further leverage geometric properties for 3D object detection.
+2.1 Query-based Object Detection Framework
+DETR [9] introduces transformers into 2D object detection, eliminating hand-designed components like non-maximum suppression, but suffering from slow convergence.
+To address this, Deformable DETR [60] proposes deformable cross-attention to accelerate training.
+DAB-DETR [28] enhances query representation by modeling queries as anchor boxes, while DN-DETR [19] stabilizes training with query denoising and auxiliary supervision.
+DETR-like approaches have been extended to 3D object detection [43, 29, 30, 33, 42, 16, 27, 5, 21, 49, 46].
+PETR [29] defines object query features and positions using learnable parameters, leveraging position embeddings within multi-head attention for feature refinement.
+DETR3D [43] and BEVFormer [24] project BEV queries onto image planes, refining them with bilinear interpolation or deformable cross-attention.
+CenterFormer [58] samples features from high-scoring heatmap keypoints for query initialization, while TransFusion [1] enhances sampled features with category embeddings.
+While extensive research has explored query initialization, DINO [54] notes that direct feature sampling in methods [58, 1, 51, 55, 44] limits performance due to inaccurate initial feature representations.
+Furthermore, despite advancements in attention-based query refinement, these approaches insufficiently leverage the geometric structure of object queries, limiting 3D spatial understanding.
+2.2 Implicit Multi-modal Fusion in Transformer
+Implicit multi-modal fusion integrates each sensor’s data through cross-attention with object queries in the transformer decoder, without relying on BEV representation for feature alignment.
+FUTR3D [5] introduces a modality-agnostic feature sampler that aggregates features from different sensors using deformable cross-attention [60].
+DeepInteraction [49] and DeepInteraction++ [50] preserve modality-specific information by using a modality interaction strategy.
+Meanwhile, TransFusion [1] updates object queries in a sequential manner by applying cross-attention separately to LiDAR and camera features.
+And CMT [46] constructs input tokens by adding 3D position embeddings to each sensor’s data before concatenating them.
+These implicit fusion methods rely on cross-attention between object queries and sensor features within the transformer decoder, providing a flexible and generalizable framework for multi-modal fusion.
+However, because it depends entirely on query-driven attention mechanisms, it struggles to effectively extract complementary features across different sensors.
+Furthermore, the iterative execution of cross-attention over the entire sensor data incurs high computational overhead and memory inefficiency.
+2.3 Explicit Multi-modal Fusion in BEV Space
+Explicit multi-modal fusion focuses on view transformation of 2D image features into BEV space [21, 31, 25, 1, 3, 35, 34, 15, 14, 22, 23, 24, 48].
+For view transformation, depth-based methods predict pixel-wise depth distributions to lift multi-view image features into BEV space [34, 15, 14, 22, 21, 31, 25].
+While these methods effectively incorporate spatial priors through depth estimation, their heavy reliance on the depth estimators significantly limits their overall robustness.
+In contrast, query-based view transformation methods [24, 48, 3, 12] project predefined 3D points onto image planes and extract features using deformable cross-attention [60].
+This approach eliminates the need for depth estimation; however, predefined positions of sampling points fail to align accurately with regions where objects are located, leading to suboptimal feature representation.
+Furthermore, it suffers from high computational overhead due to the extensive use of multi-layer transformers and ray-directional misalignment due to insufficient geometric guidance.
+Aforementioned challenges cause the inherent difficulty in establishing precise correspondences between 2D and 3D spaces, which is crucial for effective BEV representation.
+Existing approaches often suffer from feature misalignment due to depth estimation errors and also struggle with maintaining spatial consistency in transformer-based methods, while their high computational cost significantly limits real-time deployment.
+Therefore, a more efficient and geometrically grounded approach is essential for achieving accurate and robust multi-modal BEV representations.
 3 Methodology
-The proposed EVT is a novel method for 3D object detection with effective multi-modal fusion and improved transformer-based detection framework.
-The overall pipeline is shown in Fig.
+The overall pipeline of EVT is illustrated in Fig.
 2.
-First, $N_{\text{s}}$-scale perspective-view image features $\{\mathbf{PV}_{j}\mid j\in\{1,2,\dots,N_{\text{s}}\}\}$ and bird’s-eye-view (BEV) LiDAR features $\mathbf{BEV}_{\text{lidar}}\in\mathbb{R}^{C\times H\times W}$ are extracted from individual backbones, where $C$ is the feature dimension, and $H\times W$ is the size of the BEV feature map.
-Next, the proposed adaptive sampling and adaptive projection (ASAP) module fuses the multi-modal features in BEV space (Sec.
+First, $N_{\text{s}}$-scale perspective-view image features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ and BEV LiDAR features $\mathbf{BEV}_{\text{lidar}}\in\mathbb{R}^{C\times H\times W}$ are extracted from separate backbone networks, where $C$ is the feature dimension and $H\times W$ is the size of the BEV feature map.
+To fuse 2D image and LiDAR features in BEV space, the Adaptive Sampling and Adaptive Projection (ASAP) module transforms 2D image features into BEV space and then fuses them with $\mathbf{BEV}_{\text{lidar}}$ (Sec.
 3.1).
-Group-wise queries are then initialized in BEV space using the heatmap head (Sec.
-3.2), and updated in the transformer decoder based on the geometric properties of the 3D bounding box (Sec.
+For query-based 3D detection with multi-modal BEV features, group-wise mixed query selection initializes object queries based on heatmap-guided locations (Sec.
+3.2), and geometry-aware cross-attention refines them in the transformer decoder to predict 3D bounding boxes (Sec.
 3.3).
 3.1 Adaptive Sampling and Adaptive Projection
-The proposed ASAP module effectively leverages LiDAR guidance to transform multi-scale perspective-view image features into BEV space. The structure of this module is shown in Fig. 3.
-Adaptive Sampling
-The adaptive sampling (AS) module samples multi-scale image features into BEV space by predicting the height of each sampling point on the BEV plane and assigning weights to the predicted heights and image scales.
-First, a set of heights for each grid cell is generated from LiDAR features:
-$\{Z_{i}\}_{i=1}^{N_{\text{h}}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v),$ (1)
-where $(u,v)$ denotes the coordinates of a grid cell in BEV space, and $N_{\text{h}}$ denotes the number of generated heights.
-Consequently, the 3D adaptive sampling points $P=\{(X,Y,Z_{i})\}_{i=1}^{N_{\text{h}}}$ are defined using real-world coordinates $(X,Y)$ corresponding to the grid cell at $(u,v)$ and the set of heights $\{Z_{i}\}$.
-Next, the generated points $P$ are projected onto the image plane.
-Each projected point samples the multi-scale features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ across $N_{\text{s}}$ different image scales with downsampled strides $\{S_{j}\}_{j=1}^{N_{\text{s}}}$.
-Consequently, $N_{\text{h}}\times N_{\text{s}}$ projected points $(x^{j}_{i},y^{j}_{i})$ and sampled features $f^{j}_{i}$ are obtained from the projection:
-$(x^{j}_{i},y^{j}_{i})=\frac{{\rm Proj}(X,Y,Z_{i})}{S_{j}}$ (2)
-$f^{j}_{i}={\rm B}(\mathbf{PV}_{j},(x^{j}_{i},y^{j}_{i}))\in\mathbb{R}^{C},$ (3)
-where ${\rm Proj}(\cdot)$ denotes the projection of 3D points onto the image plane, and ${\rm B}(\cdot)$ denotes bilinear interpolation.
-To aggregate all the $N_{\text{s}}\times N_{\text{h}}$ sampled features $\{f^{j}_{i}\}$, adaptive sampling weights $W_{\text{as}}\in\mathbb{R}^{N_{\text{s}}\times N_{\text{h}}}$ are obtained from $\mathbf{BEV}_{\text{lidar}}$.
-$W_{\text{as}}$ determines the importance of the heights and image feature scales for each grid cell.
-The multi-scale image features in BEV space are obtained as follows:
-$W_{\text{as}}=\sigma({\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v))$ (4)
-$\mathbf{BEV}_{\text{as}}(u,v)=\sum_{j=1}^{N_{\text{s}}}\sum_{i=1}^{N_{\text{h}}}W_{\text{as}}(j,i)\cdot f^{j}_{i},$ (5)
-where $\sigma(\cdot)$ denotes the softmax function applied to all $N_{\text{s}}\times N_{\text{h}}$ elements.
-$\mathbf{BEV}_{\text{as}}$ denotes the image BEV feature map, representing the features of the ROIs from the images in BEV space by leveraging LiDAR guidance.
+The proposed ASAP module consists of two stages: Adaptive Sampling (AS) and Adaptive Projection (AP).
+In the first stage, AS selectively extracts and aggregates multi-scale perspective-view image features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ into an initial BEV representation $\mathbf{BEV}_{\text{as}}$.
+In the second stage, AP refines $\mathbf{BEV}_{\text{as}}$ to obtain the final image BEV feature map $\mathbf{BEV}_{\text{camera}}$.
+The structure of this module is shown in Fig.
+3.
 Figure 3: Overview of AS and AP.
@@
 In AP, the image BEV feature map produced by AS is refined channel-wise using the adaptive kernels generated from the LiDAR features.
+Adaptive Sampling
+To transform multi-scale image features into BEV space, the Adaptive Sampling (AS) module predicts optimal sampling heights for each BEV grid cell using LiDAR features, allowing effective feature extraction from high-relevance areas in the image (see Fig.
+4).
+First, multiple height values for each grid cell are generated from LiDAR features:
+$\{Z_{i}\}_{i=1}^{N_{\text{h}}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v),$ (1)
+where $(u,v)$ denotes the coordinates of a grid cell in BEV space, and $N_{\text{h}}$ denotes the number of generated heights.
+Consequently, the 3D sampling points $P=\{(X,Y,Z_{i})\}_{i=1}^{N_{\text{h}}}$ are defined based on real-world coordinates $(X,Y)$ corresponding to the grid cell at $(u,v)$ and the set of heights $\{Z_{i}\}$.
+Next, the generated points $P$ are projected onto the image plane.
+Each projected point samples the multi-scale features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ across $N_{\text{s}}$ different image scales with downsampled strides $\{S_{j}\}_{j=1}^{N_{\text{s}}}$.
+Consequently, $N_{\text{h}}\times N_{\text{s}}$ projected points $(x^{j}_{i},y^{j}_{i})$ and sampled features $f^{j}_{i}$ are obtained via projection:
+$(x^{j}_{i},y^{j}_{i})=\frac{{\rm Proj}(X,Y,Z_{i})}{S_{j}}$ (2)
+$f^{j}_{i}={\rm B}(\mathbf{PV}_{j},(x^{j}_{i},y^{j}_{i}))\in\mathbb{R}^{C},$ (3)
+where ${\rm Proj}(\cdot)$ denotes the projection of 3D points onto the image plane, and ${\rm B}(\cdot)$ denotes bilinear interpolation.
+To aggregate the $N_{\text{s}}\times N_{\text{h}}$ sampled features $\{f^{j}_{i}\}$, adaptive sampling weights $W_{\text{as}}\in\mathbb{R}^{N_{\text{s}}\times N_{\text{h}}}$ are derived from $\mathbf{BEV}_{\text{lidar}}$.
+$W_{\text{as}}$ determines the importance of the heights and image feature scales for each grid cell.
+The multi-scale image features in BEV space are obtained as follows:
+$W_{\text{as}}=\sigma({\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v))$ (4)
+$\mathbf{BEV}_{\text{as}}(u,v)=\sum_{j=1}^{N_{\text{s}}}\sum_{i=1}^{N_{\text{h}}}W_{\text{as}}(j,i)\cdot f^{j}_{i}\in\mathbb{R}^{C},$ (5)
+where $\sigma(\cdot)$ denotes the softmax function applied to all $N_{\text{s}}\times N_{\text{h}}$ elements. $\mathbf{BEV}_{\text{as}}$ denotes the image BEV feature map.
 Adaptive Projection
-The adaptive projection (AP) module refines the BEV feature $\mathbf{BEV}_{\text{as}}$ produced by the AS module using adaptive kernels for each grid cell.
+The Adaptive Projection (AP) module refines the BEV feature map $\mathbf{BEV}_{\text{as}}$, produced by the AS module, by applying adaptive kernels to each grid cell.
 The overall process is represented by the following equations:
@@
 $\mathbf{BEV}_{\text{camera}}(u,v)=\mathbf{BEV}_{\text{as}}(u,v)\times K_{\text{ap}}\in\mathbb{R}^{C}.$ (7)
-First, an adaptive kernel $K_{\text{ap}}\in\mathbb{R}^{C\times C}$ is generated from $\mathbf{BEV}_{\text{lidar}}$.
-This kernel is generated for each BEV grid cell to enable detailed feature representation.
-Subsequently, $\mathbf{BEV}_{\text{as}}$ is refined by applying a channel-wise linear projection using $K_{\text{ap}}$ to obtain the image BEV feature map $\mathbf{BEV}_{\text{camera}}$.
-This LiDAR-guided refinement improves feature representation and addresses the ray-directional misalignment inherent in 3D-to-2D projection methods, which suffer from occlusions or empty 3D spaces.
+First, an adaptive kernel $K_{\text{ap}}\in\mathbb{R}^{C\times C}$ is derived from $\mathbf{BEV}_{\text{lidar}}$ for each BEV grid cell.
+Then, $\mathbf{BEV}_{\text{as}}$ is refined by applying a channel-wise linear projection using $K_{\text{ap}}$ to obtain the image BEV feature map $\mathbf{BEV}_{\text{camera}}$.
+Unlike static transformations, our LiDAR-guided feature refinement leverages spatial information to effectively mitigate ray-directional misalignment, which primarily results from occlusions and empty 3D spaces (see Fig.
+5).
 Multi-modal Fusion in BEV Space
-The multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}\in\mathbb{R}^{C\times H\times W}$ is obtained by channel-wise concatenation of the image and LiDAR BEV feature maps, followed by a convolution operation:
+The multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}\in\mathbb{R}^{C\times H\times W}$ is obtained by concatenating the image and LiDAR BEV feature maps ($\mathbf{BEV}_{\text{camera}}$ and $\mathbf{BEV}_{\text{lidar}}$) along the channel dimension, followed by a convolution operation:
 $\mathbf{BEV}_{\text{fuse}}={\rm Conv}({\rm Concat}(\mathbf{BEV}_{\text{camera}},\mathbf{BEV}_{\text{lidar}})),$ (8)
-where ${\rm Concat}(\cdot)$ denotes the channel-wise concatenation.
+where ${\rm Concat}(\cdot)$ denotes channel-wise concatenation.
+3.2 Group-wise Mixed Query Selection
+Our proposed group-wise mixed query selection generates object queries for transformer-based detection frameworks.
+First, group-wise heatmaps are predicted from the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$ (from Sec.
+3.1), where each group consists of similarly sized object classes.
+The predicted heatmaps have scores ranging from 0 to 1, representing the likelihood that each BEV pixel corresponds to the center of an object.
+The heatmap head is supervised by 2D Gaussian distributions centered at each object’s location.
+Next, the top-k keypoints are selected from each heatmap group, and their positions are used as reference points for queries in BEV space.
+Inspired by DINO [54], we initialize query features solely with group-wise learnable parameters, unlike query positions.
+DINO defines query features as instance-wise learnable parameters without using any categorical priors.
+In contrast, our approach allows all queries within the same group to share these parameters, effectively capturing the common properties of their group.
+Experimentally, the group-wise shared initial embeddings outperform instance-wise embeddings for object query representation.
+Moreover, our approach outperforms traditional approaches that either define both query features and positions as learnable parameters [29, 30, 27, 46, 12] or obtain both from heatmap keypoints [58, 1, 51, 55, 44].
+Further details and analysis of our method are provided in Sec.
+4.4.
+3.3 Geometry-aware Cross-Attention
+To enhance query representations in transformer decoders, we refine deformable cross-attention [60] with corner-aware sampling for improved feature sampling and position-aware feature mixing for better feature aggregation.
+Corner-aware Sampling
+The conventional deformable cross-attention samples features around the centers of object queries.
+However, this approach struggles with objects of varying sizes and fails to capture fine-grained boundary details and spatial extent effectively.
+To address this limitation, corner-aware sampling explicitly incorporates object geometry, ensuring precise spatial alignment of sampling points with the object’s true structure.
+First, the initial sampling offsets $\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\}$ are generated from the query feature $\mathbf{q}$ using a linear layer:
+$\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\mid i\in 0,1,...,N_{\text{p}}-1\}={\rm Linear}(\mathbf{q}),$ (9)
+where $N_{\text{p}}$ denotes the number of sampling points.
+Next, the final sampling offsets $\{(\Delta x_{i},\Delta y_{i})\}$ and sampling points $\{(x_{i},y_{i})\}$ are determined via a geometric transformation, which relocates sampling points to object corners and aligns them with the object’s heading, as follows:
+$\left[\!\!\begin{array}[]{c}\Delta x_{i}\\[5.0pt] \Delta y_{i}\\ \end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}\cos\theta\!\!\!&-\!\sin\theta\!\!\\[5.0pt] \sin\theta\!\!\!&\cos\theta\!\!\!\!\\ \end{array}\!\!\right]\!\!\left[\!\!\begin{array}[]{c}{\rm I}_{j}\cdot\frac{l}{2}+\Delta x_{i}^{\prime}\\[5.0pt] {\rm I}_{j}^{\prime}\cdot\!\frac{w}{2}\!+\Delta y_{i}^{\prime}\\ \end{array}\!\!\right]\!$ (10)
+$\left[\!\!\begin{array}[]{cc}x_{i}&y_{i}\!\end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}x_{c}&y_{c}\end{array}\!\!\right]\!+\!\left[\!\begin{array}[]{cc}\!\Delta x_{i}&\Delta y_{i}\!\!\\ \end{array}\!\right],$ (11)
+where $(I_{j},I^{\prime}_{j})\,\in\{(1,1),(1,-1),(-1,1),(-1,-1)\}$.
+$j$ denotes the index of each corner and is the remainder of $i$ divided by four, the number of corners.
+The coordinates $(x_{c},y_{c})$ denote the object’s center.
+$l$, $w$, and $\theta$ denote the length, width, and yaw of the predicted bounding box, respectively, obtained from the regression head of the previous transformer layer.
+In the first decoder layer, we initialize $l$, $w$, $\theta$ to zero.
+The features are sampled using bilinear interpolation at each sampling point on the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$, as follows:
+$g_{i}={\rm B}(\mathbf{BEV}_{\text{fuse}},(x_{i},y_{i}))\in\mathbb{R}^{C},$ (12)
+where ${\rm B}(\cdot)$ denotes the bilinear interpolation, and $g_{i}$ represents the sampled features at $(x_{i},y_{i})$.
+Position-aware Feature Mixing
+Given the sampled features from different corners, the key challenge is how to decode them while maintaining their spatial relationships.
+While AdaMixer [10] is designed for feature decoding, it struggles with the geometric transformations introduced during corner-aware sampling, which complicate the association between sampled features and their transformed sampling locations (see Tab.
+4 (e)).
+To address this, we propose position-aware feature mixing, which incorporates positional embeddings that encode the transformed sampling offsets, allowing for more structured feature aggregation.
+First, the sampling offsets are embedded as position vectors using sinusoidal position encoding [38], followed by a linear layer.
+Then, the position-aware sampled features $G\in\mathbb{R}^{N_{\text{p}}\times C}$ are obtained by element-wise addition of the sampled feature $g_{i}$ and the position vector $e_{i}$:
+$e_{i}$ $={\rm Linear}(\Phi_{pos}((x_{i},y_{i})))\in\mathbb{R}^{C}$ (13)
+$G$ =$=\{(g_{i}+e_{i})\mid i\in{0,1,...,N_{\text{p}}-1}\},$ (14)
+where $\Phi_{pos}(\cdot)$ denotes sinusoidal position encoding.
+Subsequently, adaptive channel mixing is applied to $G$ using the dynamic weights $W_{c}$ generated from the query feature $\mathbf{q}$ to obtain the channel-mixed feature $G_{c}$:
+$W_{c}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{C\times C}$ (15)
+$G_{c}$ $={\rm ReLU}({\rm LN}(G\times W_{c})).$ (16)
+Next, adaptive spatial mixing is applied to the spatial dimensions of $G_{c}$ using dynamic weights $W_{s}$ to obtain the spatial-mixed feature $G_{cs}$:
+$W_{s}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{N_{\text{p}}\times N_{\text{p}}}$ (17)
+$G_{cs}$ $={\rm ReLU}({\rm LN}(G_{c}^{T}\times W_{s})).$ (18)
+Finally, the query feature is formulated as follows:
+$\mathbf{q^{\prime}}=\mathbf{q}+{\rm Linear}({\rm Flatten}(G_{cs})),$ (19)
+where $\mathbf{q^{\prime}}$ represents the refined query feature.
+By integrating positional embeddings at the feature level, this formulation enhances spatial awareness in query representation, ensuring more robust geometric reasoning in object queries.
+4 Experiments
 Method Modality NDS (val) mAP (val) NDS (test) mAP (test)
@@
 UniTR [40] LC 73.3 70.5 74.5 70.9
+FusionFormer [12] LCT 74.1 71.4 75.1 72.6
 EVT (Ours) LC 74.6 72.1 75.3 72.6
-Table 1: Performance comparison on the nuScenes validation and test sets. The results are obtained without model ensemble or test-time augmentation. ‘L’ and ‘C’ denote LiDAR and Camera, respectively.
-3.2 Group-wise Query Initialization
-The proposed group-wise query initialization method effectively generates queries for transformer-based detection frameworks.
-First, group-wise heatmaps are predicted from the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$.
-The groups are predefined, with each group containing similarly-sized classes.
-The predicted heatmaps have scores ranging from 0 to 1, representing the likelihood that each BEV pixel corresponds to the center of an object.
-The heatmap head is supervised by 2D Gaussian distributions centered at each object’s location.
-Next, the top-k keypoints are selected from each heatmap group.
-The positions of these keypoints are used as reference points for queries in BEV space.
-This method initializes each query feature using group-wise learnable parameters.
-Queries within the same group share these parameters, which enables them to effectively capture the characteristics of group.
-This demonstrates better performance in the multi-layer transformer decoder than traditional query initialization approaches that sample features directly from the input feature map [47, 1].
-Further details are provided in Sec.
-4.4.
-3.3 Enhanced Query Update Framework
-For more effective query updates within a transformer decoder based on deformable attention [49], we propose the corner-aware sampling and position-embedded feature mixing methods.
-Corner-aware Sampling.
-The proposed corner-aware sampling method shifts the sampling reference from the center of the object to each corner and rotates sampling offsets according to the object’s heading.
-This approach addresses the limitations of existing center-based sampling methods, which generate sampling offsets around the object’s center, resulting in difficulties capturing the full spatial extent and shape of the objects.
-First, the initial points $\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\}$ are generated from the query feature $\mathbf{q}$ using a linear layer:
-$\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\mid i\in 0,1,...,N_{\text{p}}-1\}={\rm Linear}(\mathbf{q}),$ (9)
-where $N_{\text{p}}$ denotes the number of sampling points.
-Subsequently, the sampling offsets $\{(\Delta x_{i},\Delta y_{i})\}$ and sampling points $\{(x_{i},y_{i})\}$ are determined via a geometric transformation as follows:
-$\left[\!\!\begin{array}[]{c}\Delta x_{i}\\[5.0pt] \Delta y_{i}\\ \end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}\cos\theta\!\!\!&-\!\sin\theta\!\!\\[5.0pt] \sin\theta\!\!\!&\cos\theta\!\!\!\!\\ \end{array}\!\!\right]\!\!\left[\!\!\begin{array}[]{c}{\rm I}_{j}\cdot\frac{l}{2}+\Delta x_{i}^{\prime}\\[5.0pt] {\rm I}_{j}^{\prime}\cdot\!\frac{w}{2}\!+\Delta y_{i}^{\prime}\\ \end{array}\!\!\right]\!$ (10)
-$\left[\!\!\begin{array}[]{cc}x_{i}&y_{i}\!\end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}x_{c}&y_{c}\end{array}\!\!\right]\!+\!\left[\!\begin{array}[]{cc}\!\Delta x_{i}&\Delta y_{i}\!\!\\ \end{array}\!\right],$ (11)
-where $(I_{j},I^{\prime}_{j})\,\in\{(1,1),(1,-1),(-1,1),(-1,-1)\}$.
-$j$ denotes the index of each corner and is the remainder of $i$ divided by four, the number of corners.
-$(x_{c},y_{c})$ denote the coordinates of the object’s center.
-$l$, $w$, and $\theta$ denote the length, width, and yaw of the predicted bounding box, respectively, obtained from the regression head of the previous transformer decoder layer.
-In the first decoder layer, where there is no preceding layer to provide bounding boxes, the equation is satisfied by setting $l$, $w$, and $\theta$ to zero.
-The features are sampled by bilinear interpolation at each sampling point on the multi-modal fused features $\mathbf{BEV}_{\text{fuse}}$ as follows:
-$g_{i}={\rm B}(\mathbf{BEV}_{\text{fuse}},(x_{i},y_{i}))\in\mathbb{R}^{C},$ (12)
-where ${\rm B}(\cdot)$ denotes the bilinear interpolation. $g_{i}$ represents the sampled feature that captures important key tokens by leveraging the geometric properties of the query’s bounding box.
-Position-embedded Feature Mixing.
-Accurate locations of the sampled features are essential for providing a more detailed representation of the bounding boxes.
-In AdaMixer [9], the adaptive channel and spatial mixing are combined to leverage query guidance and effectively decode sampled features.
-However, this method is less effective for decoding the outputs of the corner-aware sampling, as shown in Tab.
-3 (e), which involves complex geometric transformations of the sampling points.
-To address this, we propose position-embedded feature mixing, an enhancement based on AdaMixer that improves the decoding capability by incorporating embeddings for sampling offsets.
-First, the sampling offsets are embedded as position vectors using sinusoidal position encoding [33], followed by a linear layer.
-Then, the position-embedded sampled features $G\in\mathbb{R}^{N_{\text{p}}\times C}$ are obtained by element-wise addition of the sampled feature $g_{i}$ and the position vector $e_{i}$:
-$e_{i}$ $={\rm Linear}(\Phi_{pos}((x_{i},y_{i})))\in\mathbb{R}^{C}$ (13)
-$G$ =$=\{(g_{i}+e_{i})\mid i\in{0,1,...,N_{\text{p}}-1}\},$ (14)
-where $\Phi_{pos}(\cdot)$ denotes sinusoidal position encoding.
-Subsequently, adaptive channel mixing is applied to $G$ using the dynamic weights $W_{c}$ generated from the query feature $\mathbf{q}$ to obtain the channel-mixed feature $G_{c}$:
-$W_{c}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{C\times C}$ (15)
-$G_{c}$ $={\rm ReLU}({\rm LN}(G\times W_{c})).$ (16)
-Next, adaptive spatial mixing is applied to the spatial dimensions of $G_{c}$ using dynamic weights $W_{s}$ to obtain the spatial-mixed feature $G_{cs}$:
-$W_{s}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{N_{\text{p}}\times N_{\text{p}}}$ (17)
-$G_{cs}$ $={\rm ReLU}({\rm LN}(G_{c}^{T}\times W_{s})).$ (18)
-Finally, the query feature is updated as follows:
-$\mathbf{q^{\prime}}=\mathbf{q}+{\rm Linear}({\rm Flatten}(G_{cs})),$ (19)
-where $\mathbf{q^{\prime}}$ represents the updated query feature that is enhanced for a more accurate feature representation by incorporating the positional information of each sampled point.
-4 Experiments
-Figure 4: Visualization of projected sampling points for each object.
-The top row shows the projections of predefined 3D points, and the bottom row shows the projections of the points generated by the AS module.
-Points of each object are denoted by different colors.
-Figure 5: Comparison of BEV feature maps from different methods: (a) Vanilla, (b) AS, (c) AP, and (d) ASAP.
-Vanilla shows less informative and unaligned features.
-AS improves feature representation using adaptive sampling.
-AP corrects ray-directional misalignment.
-ASAP integrates both AS and AP, leading to the most refined and well-aligned feature representation.
+Table 1: Performance comparison on the nuScenes validation and test sets.
+The results are obtained without model ensemble or test-time augmentation.
+‘L’, ‘C’ and ‘T’ denote LiDAR, camera and temporal fusion, respectively.
 4.1 Implementation Details
-For the image backbone, either ResNet-50 [10] with a resolution of $704\times 256$ or V2-99 [15] with $1600\times 640$ is used with the feature pyramid network (FPN) [22].
-For the LiDAR backbone, VoxelNet [46] is used with an ROI of $[-54.0m,54.0m]$ along $(X,Y)$ and $[-5.0m,3.0m]$ along $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$.
-In ASAP, we use 4 points for each grid cell and the last two stages of each image backbone for view transformation.
-The size of the unified BEV feature map size is $180\times 180$.
-The object categories are divided into 6 groups, with 150 queries per group, resulting in a total of 900 queries.
-The transformer decoder has 6 layers, and the corner-aware sampling generates 16 sampling points per query.
-The feature dimension is set to 256.
+For the image backbone, we use ResNet-50 [11] with a resolution of $704\times 256$ or V2-99 [18] with $1600\times 640$ with FPN [26].
+The LiDAR backbone is VoxelNet [57] with an ROI of $[-54.0m,54.0m]$ in $(X,Y)$ and $[-5.0m,3.0m]$ in $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$.
+In ASAP, we use four sampling points for each grid cell.
+The multi-modal BEV feature map size is $180\times 180$.
+Following [52], object groups in 3.2 are defined as: (1) car, (2) truck, construction vehicle, (3) bus, trailer, (4) barrier, (5) motorcycle, bicycle, (6) pedestrian, traffic cone.
+Each group contains 150 queries, resulting in a total of 900 queries, and the corner-aware sampling generates 16 sampling points per query, both of which were empirically determined.
+The transformer decoder has six layers, and the feature dimension is set to 256.
 Our model is trained on 8 RTX 3090 GPUs with a batch size of 16.
@@
 No model ensemble or test-time augmentation is applied during inference.
+Figure 4: Visualization of projected sampling points for each object.
+The top row shows the projections of predefined 3D points, and the bottom row shows the projections of the points generated by the AS module.
+Points of each object are denoted by different colors.
+(a) vanilla
+(b) only AS
+(c) only AP
+(d) ASAP
+Figure 5: Comparison of BEV feature maps from different methods: (a) Vanilla, (b) AS, (c) AP, and (d) ASAP.
+Vanilla shows less informative and unaligned features.
+AS improves feature representation using adaptive sampling.
+AP corrects ray-directional misalignment.
+ASAP integrates both AS and AP, leading to the most refined and well-aligned feature representation.
 4.2 Dataset and Metric
@@
 1, we compare EVT and its LiDAR-only model, EVT-L, with existing methods on the nuScenes validation and test sets.
-The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, outperforming all existing state-of-the-art approaches on both the nuScenes validation and test sets.
+The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, surpassing all previous approaches on both the nuScenes validation and test sets.
 In particular, it surpasses recent methods, such as UniTR [40] by 0.8% NDS and 1.7% mAP, MSMDFusion [17] by 1.3% NDS and 1.1% mAP, and SparseFusion [45] by 1.5% NDS and 0.6% mAP.
@@
 These results validate the effectiveness of the proposed query initialization and cross-attention mechanisms in enhancing 3D object detection.
-LiDAR Camera AS AP NDS mAP
-(a) 71.7 66.4
-(b) 72.7 69.1
-(c) 73.5 70.6
-(d) 73.7 70.7
+LiDAR Camera AS AP NDS mAP FPS
+(a) 71.7 66.4 12.1
+(b) 72.7 69.1 8.5
+(c) 73.5 70.6 8.5
+(d) 74.1 71.1 8.3
 Table 2: Ablation study of the proposed VT method. The proposed ASAP shows a significant performance improvement compared to the vanilla VT method (b), which projects predefined 3D points.
-Figure 6: Comparison of query initialization strategies on the nuScenes validation set.
-For the multi-layer transformer decoder, group-wise query initialization without feature sampling achieves the best performance.
-Attention Reference scale rotate feature mixing position-embedded NDS mAP
+Method # Decoder Layers & NDS (%)
+1 2 3 4 5 6
+(a) Learnable Init. 56.8 64.4 67.5 68.7 68.9 69.6
+(b) Heatmap Init. 70.4 70.5 70.5 70.3 70.5 70.7
+(c) Mixed init. 69.2 70.3 70.5 70.5 70.7 71.1
++ Group-wise 69.8 71.1 71.3 71.2 71.4 71.7
+Table 3: Comparison of query initialization strategies on the nuScenes validation set.
+(a) Fully learnable initialization, (b) Fully heatmap-based initialization, (c) Our mixed initialization strategy: the first row denotes instance-wise mixed query selection and the second row denotes group-wise mixed query selection.
+Attention Reference scale rotate feature mixing position-aware NDS mAP
 (a) Standard Center 69.6 65.3
@@
 (f) 71.7 66.4
-Table 3: Comparison of the cross-attention methods within the transformer decoder. The proposed corner-aware sampling and position-embedded feature mixing achieve significant performance improvements.
-Initial Query Formulation NDS mAP
-BEV Reference Point 71.7 66.4
-Bbox from regression head 71.4 66.3
-Bbox from learnable parameters 70.9 66.0
-Table 4: Comparison of the initial query formulations for the first layer of the transformer decoder.
+Table 4: Comparison of attention methods within the transformer decoder.
+The proposed geometry-aware cross-attention, which includes corner-aware sampling and position-aware feature mixing, achieves significant performance improvements.
+Method NDS mAP
+ResNet-50 baseline (24 epochs) 47.8 37.2
++ geometry-aware cross-attention 49.1 37.8
+ResNet-50 baseline (90 epochs) 53.5 42.7
++ geometry-aware cross-attention 54.7 43.4
+Table 5: Impact of geometry-aware cross-attention on StreamPETR [42]. All models are trained in our experiments.
 4.4 Ablation Studies
 In this section, we describe the validation of each component of the proposed method on the nuScenes validation set.
-Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [48] and the query denoising strategy [16].
+Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [59] and the denoising strategy [19].
 Adaptive Sampling and Adaptive Projection
-Tab.
-2 displays the results of the ablation conducted on ASAP by retraining the entire model.
+In Tab.
+2, the ablation results for ASAP are obtained by retraining the entire model.
 In these experiments, ResNet [11] is used as the backbone network with a resolution of $704\times 256$.
@@
 4.
-The sampling points of AS are adaptively generated in the ROIs.
-In (d), the entire ASAP achieves improvements of 1.0% NDS and 1.6% mAP compared to the vanilla VT method (b).
+The sampling points of AS are adaptively generated in highly object-relevant regions of the image.
+In (d), the entire ASAP achieves improvements of 1.4% NDS and 2.0% mAP compared to (b), while maintaining high efficiency with only a 3ms latency increase.
 The BEV feature maps of each component are visualized in Fig.
@@
 ASAP effectively transforms image features into BEV space and resolves ray-directional misalignment.
-Group-wise Query Initialization
-Different query initialization methods using the heatmap head are ablated in Fig.
-6.
-All methods utilize a top-K keypoints selection approach to define queries, but differ in their specific implementations.
-(a) directly samples features from the input BEV feature map and combines them with group-wise embedding, which is similar to the approach commonly used in 3D object detectors [47, 1].
-(b) initializes queries with group-wise embedding using global top-K selection.
-(c) is our proposed group-wise query initialization strategy, which initializes queries using top-K selection and embedding for each group.
-In the results, the direct feature sampling approach in (a) is more effective in a single-layer transformer decoder, outperforming (c) by 0.6% NDS.
-However, in a multi-layer transformer decoder, this leads to performance degradation.
-Specifically, with 6 layers, (c) surpasses (a) by 1.0% NDS.
-Additionally, group-wise top-K (c) consistently outperforms global top-K (b) across all layers by approximately 0.1% NDS.
-Enhanced Query Update Framework
-We ablate the corner-aware sampling method, as demonstrated in Tab.
+Group-wise Mixed Query Selection
+As shown in Tab.
+3, we conduct an ablation study on query initialization strategies.
+(a) fully learnable initialization, where both features and positions are learnable [29, 30, 27, 46, 12].
+(b) fully heatmap-based initialization, where positions are obtained from high-score heatmap keypoints, and features are sampled at those locations [58, 1, 51, 55, 44].
+(c) our mixed initialization strategy, which combines heatmap-derived positions with either instance-wise or group-wise embeddings.
+These results highlight the importance of query initialization.
+Notably, fully learnable initialization without any prior information consistently yields the lowest performance.
+While (b) achieves the best performance in a single-layer transformer decoder, (c) outperforms all other strategies in multi-layer settings, improving NDS by over 1% at the last layer.
+Additionally, group-wise embeddings allow each group to learn more generalized feature representations, leading to meaningful improvements.
+Geometry-aware Cross-Attention
+We ablate the corner-aware sampling method, as shown in Tab.
 4 (a)-(d).
-(a) is the standard multi-head cross attention [33], and (b)-(c) are the standard deformable attention [49].
-Specifically, in (c), the sampling offsets are scaled according to the bounding box size and rotated according to the heading.
-(d) is the proposed corner-aware sampling method, which samples features from the bounding box corners of each query and rotates them according to the heading.
-The corner-aware sampling method (d) achieves improvements of 0.8% NDS and 0.5% mAP compared to method (c).
-The results of the ablation study on the feature mixing method are displayed in Tab.
-3 (e)-(f).
-In (e), the standard feature mixing method of AdaMixer [9] does not yield any improvements over our method.
-In (f), the proposed position-embedded feature mixing demonstrates improvements of 0.4% NDS and 0.7% mAP compared to (d).
+(a) employs multi-head attention [38], and (b)-(c) use deformable attention [60].
+Specifically, in (c), the sampling offsets are scaled based on the bounding box size and rotated according to the heading.
+(d), the proposed corner-aware sampling method, samples features from bounding box corners aligned with the heading, achieving improvements of 0.8% NDS and 0.5% mAP compared to (c).
+In (e) and (f), AdaMixer [10] fails to improve performance, as it does not effectively preserve the structured sampling introduced by corner-aware sampling.
+In contrast, our position-aware feature mixing explicitly incorporates positional embeddings, leading to improvements of 0.4% NDS and 0.7% mAP compared to (d).
+As a result, our proposed geometry-aware cross-attention achieves overall improvements of 1.2% NDS and 1.2% mAP.
+Additionally, as shown in Tab.
+5, we validate the applicability of geometry-aware cross-attention by integrating it into a camera-only 3D detector.
+Modifying only the cross-attention layers in StreamPETR [42] improves performance without further adjustments.
+Initial Query Formulation NDS mAP
+Bbox from regression head 71.4 66.3
+Bbox from learnable parameters 70.9 66.0
+BEV Reference Point 71.7 66.4
+Table 6: Comparison of the initial query formulations for the first layer of the transformer decoder.
 Initial Query Formulation for Transformer
 As described in Sec.
-3.3, the corner-aware sampling is applied starting from the second transformer decoder layer.
-We experimented on two different approaches to incorporate the corner-aware sampling into all transformer decoder layers.
-The first approach involves adding regression heads during the query initialization stage (Sec.
-3.2) to predict the bounding boxes, whereas the second approach uses learnable bounding boxes from SparseBEV [23].
+3.3, the corner-aware sampling method is applied starting from the second transformer layer.
+We explore two different approaches to extend the corner-aware sampling to all transformer layers.
+The first approach involves adding regression heads during query initialization in Sec.
+3.2 to predict bounding boxes, whereas the second approach uses learnable bounding boxes.
 However, as shown in Tab.
-4, neither of these approaches led to any noticeable performance improvement.
-Furthermore, the experiment using learnable bounding boxes results in performance degradation.
+6, neither approach resulted in any noticeable performance gains, and the experiment with learnable bounding boxes even led to performance degradation.
 5 Conclusion
-In this paper, we have proposed EVT, a novel transformer-based 3D object detector.
-The proposed view transformation method, ASAP, enables robust and efficient LiDAR-camera fusion by leveraging LiDAR guidance.
-In addition, the proposed group-wise query initialization method provides effective queries for a multi-layer transformer decoder.
-The corner-aware sampling and position-embedded feature mixing methods effectively utilize geometric properties to update queries, and these methods can easily be extended to other models.
+We propose EVT, a novel multi-modal 3D object detector based on BEV representation, enhancing both efficiency and accuracy.
+Our method introduces ASAP, an efficient LiDAR-camera fusion method that leverages LiDAR guidance for accurate view transformation.
+Additionally, the proposed group-wise mixed query selection improves initial feature representation through shared embeddings.
+The geometry-aware cross-attention refines queries using geometric properties and can be easily extended to other models.
 We expect EVT to provide valuable insights into multi-modal 3D object detection.